Out of 25 employees of a company, 5 are engineers. Three employees are selected at random for granting leave. What is the probability that (i) all the three are engineers?(ii) none of them is an engineer? (iii) at least one of them is an engineer? 


1
Expert's answer
2020-07-13T19:04:01-0400

Solution:

(i) Let A1 be event that all the three are engineers.

Then

P(A1)=m/n

m=5C3=5!/3!/2!=10,

n=25C3=25!/3!/22!=2300,

P(A1)=10/2300=1/230.


(ii) Let A2 be event that "none of them is an engineer".

Then

P(A2)=m/n

m=20C3=20!/3!/17!=1140,

n=25C3=25!/3!/22!=2300,

P(A2)=1140/2300=57/115.


(iii) Let A3 be event that "at least one of them is an engineer".

Then

"A_2\bigcup A_3 =U",

"A_2 \bigcap A_3 =\varnothing"


P(A2)+P(A3)=1

P(A3)=1-P(A2)

P(A3)=1-57/115=58/115.


Answer:

(i)   the probability that all the three are engineers is equal to 1/230;

(ii) the probability that none of them is an engineer is equal to 57/115;

(iii) the probability that at least one of them is an engineer is equal to 58/115.